# Subject-Consistent and Pose-Diverse Text-to-Image Generation

### v1

@@ 4.2 Experimental Results
Qualitative comparison. As shown in Figure 3, our CoDi achieves superior visual quality in terms of pose diversity, subject consistency, and prompt fidelity. Our CoDi preserves the pose diversity of Vanilla SDXL podell2023sdxl while overcoming its limitation in subject consistency. In comparison, ConsiStory tewel2024training and StoryDiffusion zhou2024storydiffusion achieve subject consistency at the cost of pose diversity. For example, in the scientist scenario, the man exhibits nearly identical body poses. Although 1-Prompt-1-Story liu2025one maintains strong layout and pose diversity in both cases, its subject consistency remains limited.
@@
Quantitative comparison. Table 1 presents a quantitative comparison. (1) Across all three subject consistency metrics—CLIP-I hessel2021clipscore , DINO-v2 oquab2023dinov2 , and DreamSim fu2023dreamsim —our CoDi achieves the best performance, demonstrating superior identity preservation across instances. In particular, our method obtains the lowest DreamSim score (0.2136), indicating closer alignment with human perceptual similarity than competing methods. (2) In terms of pose diversity, CoDi achieves the highest score (0.0758), closely matching Vanilla SDXL (0.0772). This demonstrates its ability to preserve the inherent pose diversity of the diffusion model while maintaining subject consistency. (3) For prompt fidelity, CoDi performs competitively—ranking second only to ConsiStory and comparable to Vanilla SDXL. These results demonstrate CoDi’s ability to achieve subject consistency without compromising pose diversity or prompt alignment.
@@ 4.3 Ablation Studies
Figure 4 visualizes the effective of each module. Compared to Vanilla SDXL, applying IT preserves the original pose and improves subject consistency, but some details, such as facial identity, remain suboptimal. In contrast, IR alone enhances fine-grained consistency, but leads to nearly identical poses across images, resulting in a substantial loss of diversity. As shown in the bottom row, combining IT and IR improves both coarse and fine-grained consistency without compromising pose diversity.
Table 3: Inference Time and Memory Usage. We report the inference time (in seconds) and peak GPU memory usage on a single A6000 GPU for generating a set of five images from a prompt set with a resolution of $1024\times 1024$. StoryDiffusion zhou2024storydiffusion is excluded due to excessive GPU memory consumption beyond the A6000’s limit.
Method Inference time (s) GPU memory (GB)
Vanilla SDXL podell2023sdxl 77.67 35.58
1-Prompt-1-Story liu2025one 115.85 17.13
ConsiStory tewel2024training 113.88 46.60
CoDi (ours) 154.89 45.20
@@
Inference time and memory usage. We measure the inference time and memory usage of different SCG methods on a single A6000 GPU, as shown in Table 3. We report the wall-clock time for generating a set of five images from a prompt set (since the baseline method ConsiStory tewel2024training performs cross-image attention across a batch of images) at a resolution of $1024\times 1024$. Based on Table 3, our method CoDi exhibits slightly higher inference time (154.89s) and comparable GPU memory usage (45.20GB) relative to ConsiStory. While 1-Prompt-1-Story liu2025one is the most memory-efficient, it compromises subject consistency. Note that we exclude StoryDiffusion zhou2024storydiffusion due to excessive GPU memory usage beyond the A6000’s limit at $1024\times 1024$ resolution (its original setting uses $768\times 768$).
@@ 5 Conclusion
In this paper, we propose CoDi, a novel training-free framework that addresses the trade-off between subject consistency and pose diversity. CoDi comprises two key components: identity transport (IT) and identity refinement (IR). During early denoising steps, IT aligns features across instances by optimally transporting the identity subject’s features to each, while preserving pose diversity. IR further refines subject consistency by aligning instance features with the salient attributes of the identity subject in the later denoising steps. The effectiveness of our CoDi is demonstrated by its state-of-the-art performance in achieving subject consistency and maintaining pose diversity.
@@ Appendix A Additional Implementation Details
We apply Otsu’s thresholding [25] to obtain the binary subject mask $M$:
@@ Appendix B Additional Evaluation Details
Pose diversity score. We begin by extracting normalized 2D human keypoints and their confidence scores from each target image using ViTPose [38], a SoTA transformer-based model known for its high accuracy and robustness in human pose estimation. Each image $\mathbf{x}$ is represented by a set of $H$ keypoint locations $\mathbf{p}$ and their confidences $\bm{\beta}$.
@@
where each keypoint $\mathbf{p}_{i}=(p_{i}^{x},p_{i}^{y})$ is normalized by the image width and height and $\beta_{i}\in[0,1]$ denotes its confidence score. To ensure robustness, we discard keypoints with confidence scores below a threshold $\tau$. For a pair of target images $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, we retain only the indices of keypoints that are valid in both images. We then perform Procrustes method [34] to remove global variations in translation, rotation, and scale by aligning $\mathbf{p}_{i}$ to $\mathbf{p}_{j}$. Specifically, we first compute the centroids of the keypoints which are denoted as $\bm{\mu}_{i}$ and $\bm{\mu}_{j}$. We then center both keypoint sets by subtracting their respective centroids and normalize their $\ell_{2}$ norm:
@@
$\mathbf{U},\,\boldsymbol{\Sigma},\,\mathbf{V}^{\top}=\mathrm{SVD}(\bar{\mathbf{p}}_{i}^{\top}\bar{\mathbf{p}}_{j}),\quad\mathbf{R}=\mathbf{V}^{\top}\mathbf{U}^{\top}.$ (19)
@@
$\gamma=\frac{\|\bar{\mathbf{p}}_{j}\|_{2}}{\|\bar{\mathbf{p}}_{i}\|_{2}}\cdot\mathrm{tr}(\boldsymbol{\Sigma}).$ (20)
@@
$\hat{\mathbf{p}}_{i}=\gamma\cdot\bar{\mathbf{p}}_{i}\mathbf{R}+\boldsymbol{\mu}_{j}.$ (21)
The pose diversity score between a pair of images $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ is computed as the average Euclidean distance between $\hat{\mathbf{p}}_{i}$ and $\mathbf{p}_{j}$. To analyze pose diversity under different confidence thresholds $\tau$, we compare the pose diversity scores of various methods across a range of $\tau$ values. As shown in the Fig. 6, our CoDi consistently outperforms other SCG methods under all $\tau$ settings, and achieves performance comparable to Vanilla SDXL [26]. We use $\tau=0.7$ in our experiments to balance keypoint reliability and coverage.
Figure 6: Pose diversity scores across different confidence thresholds $\tau$. Our CoDi  consistently outperforms other SCG methods and performs comparably to Vanilla SDXL [26].
@@ Appendix C Limitations
Similar to prior subject-mask-based methods such as ConsiStory [35], our CoDi  framework relies on cross-attention scores to extract subject masks and estimate image token importance in the OT Plan. Occasionally, the pre-trained diffusion model assigns higher attention to background regions than to the subject, as shown in Fig. 7, which hinders the effective transport of identity features $\mathbf{X}_{\mathsf{id}}$ from the reference image $\mathbf{x}_{\mathsf{id}}$ to target image $\mathbf{x}_{n}$, resulting in subject inconsistency. However,such failures are rare in practice (under 5%) and can be solved by simply changing the seed.
Appendix D Additional Results

### v2

@@ 4.2 Experimental Results
Qualitative comparison. As shown in Figure 3, our CoDi achieves superior visual quality in terms of pose diversity, subject consistency, and prompt fidelity. Our CoDi preserves the pose diversity of Vanilla SDXL Podell et al. (2023) while overcoming its limitation in subject consistency. In comparison, ConsiStory Tewel et al. (2024) and StoryDiffusion Zhou et al. (2024) achieve subject consistency at the cost of pose diversity. For example, in the scientist scenario, the man exhibits nearly identical body poses. Although 1Prompt1Story Liu et al. (2025) maintains strong layout and pose diversity in both cases, its subject consistency remains limited.
@@
Quantitative comparison. Table 1 presents a quantitative comparison. (1) Across all three subject consistency metrics—CLIP-I Hessel et al. (2021), DINO-v2 Oquab et al. (2023), and DreamSim Fu et al. (2023)—our CoDi achieves the best performance, demonstrating superior identity preservation across instances. In particular, our method obtains the lowest DreamSim score (0.2136), indicating closer alignment with human perceptual similarity than competing methods. (2) In terms of pose diversity, CoDi achieves the highest score (0.0758), closely matching Vanilla SDXL (0.0772). This demonstrates its ability to preserve the inherent pose diversity of the diffusion model while maintaining subject consistency. (3) For prompt fidelity, CoDi performs competitively—ranking second only to ConsiStory and comparable to Vanilla SDXL. These results demonstrate CoDi’s ability to achieve subject consistency without compromising pose diversity or prompt alignment.
@@ 4.3 Ablation Studies
Figure 4 visualizes the effect of each module. Compared to Vanilla SDXL, applying IT preserves the original pose and improves subject consistency, but some details, such as facial identity, remain suboptimal. In contrast, IR alone enhances fine-grained consistency, but produces nearly identical poses across images, substantially reducing diversity. As shown in the bottom row, combining IT and IR improves both coarse and fine-grained consistency without compromising pose diversity.
@@
4.4 Multi-subject Generation
Both IT and IR are independently applied to each subject and are easily extendable, enabling our CoDi to naturally support multi-subject generation. For each subject, we perform IT and extract its most salient features for IR , which effectively prevents feature interference across subjects and enhances subject consistency. As shown in Fig. 6, our CoDi preserves multi-subject consistency while maintaining their pose diversity.
Figure 6: Multi-subject generation. CoDi consistently preserves identities of multiple subjects while maintaining diverse poses and spatial layouts.
4.5 Different Style Generation
CoDi first transports the identity features from the reference image during the IT stage, and in the IR stage, the diffusion model refines the subject with a specific style. As shown in Fig. 7, CoDi generates images with consistent subject appearance and diverse styles.
Figure 7: CoDi can generate images with consistent subject appearance and different style.
4.6 User Study
We conducted a user study to compare our method with state-of-the-art approaches. A total of 30 prompt sets were randomly sampled, each consisting of four fixed-length prompts. Thirty-nine participants were asked to evaluate which method demonstrated the best overall performance of the generated images in terms of subject consistency, pose/layout diversity, and prompt fidelity. As shown in Table 3, CoDi achieved the highest overall preference, surpassing the second-best method (StoryDiffusion) by 24.01%.
Table 3: User study with 39 participants evaluating T2I SCG methods based on human preference.
Method ConsiStory StoryDiffusion 1Prompt1Story CoDi (ours)
Percent (%) $\uparrow$ 19.06 21.20 14.53 45.21
@@ 5 Conclusion
In this paper, we propose CoDi, a novel training-free framework that addresses the trade-off between subject consistency and pose diversity. CoDi comprises two key components: identity transport (IT) and identity refinement (IR). During early denoising steps, IT aligns features across instances by optimally transporting the identity subject’s features, while preserving pose diversity. IR further refines subject consistency by aligning instance features with the salient attributes of the identity subject in the later denoising steps. The effectiveness of our CoDi is demonstrated by its state-of-the-art performance in achieving subject consistency and maintaining pose diversity.
Acknowledgments
This work was supported by the Gusu Innovation and Entrepreneur Leading Talents: No. ZXL2024362, Natural Science Foundation of Jiangsu Province: BK20241198, and Natural Science Foundation of China: No. 62406135.
@@ Appendix A Additional Implementation Details
We apply Otsu’s thresholding Otsu et al. (1975) to obtain the binary subject mask $M$:
@@
Figure 8: Pose diversity scores across different confidence thresholds $\tau$. Our CoDi  consistently outperforms other SCG methods by a clear margin under all $\tau$ settings.
@@ Appendix B Additional Evaluation Details
Pose diversity score. We begin by extracting normalized 2D human keypoints and their confidence scores from each target image using ViTPose Xu et al. (2022), a SoTA transformer-based model known for its high accuracy and robustness in human pose estimation. Each image $\mathbf{x}$ is represented by a set of $H$ keypoint locations $\mathbf{p}$ and their confidences $\bm{\beta}$.
@@
where each keypoint $\mathbf{p}_{i}=(p_{i}^{x},p_{i}^{y})$ is normalized by the image width and height and $\beta_{i}\in[0,1]$ denotes its confidence score. To ensure robustness, we discard keypoints with confidence scores below a threshold $\tau$. For a pair of target images $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, we retain only the indices of keypoints that are valid in both images. We then perform Procrustes method Schönemann (1966) to remove global variations in translation, rotation, and scale by aligning $\mathbf{p}_{i}$ to $\mathbf{p}_{j}$. Specifically, we first compute the centroids of the keypoints which are denoted as $\bm{\mu}_{i}$ and $\bm{\mu}_{j}$. We then center both keypoint sets by subtracting their respective centroids and normalize their $\ell_{2}$ norm:
@@
$\mathbf{U},\,\bm{\Sigma},\,\mathbf{V}^{\top}=\mathrm{SVD}(\bar{\mathbf{p}}_{i}^{\top}\bar{\mathbf{p}}_{j}),\quad\mathbf{R}=\mathbf{V}^{\top}\mathbf{U}^{\top}.$ (19)
@@
$\gamma=\frac{\|\bar{\mathbf{p}}_{j}\|_{2}}{\|\bar{\mathbf{p}}_{i}\|_{2}}\cdot\mathrm{tr}(\bm{\Sigma}).$ (20)
@@
$\hat{\mathbf{p}}_{i}=\gamma\cdot\bar{\mathbf{p}}_{i}\mathbf{R}+\bm{\mu}_{j}.$ (21)
The pose diversity score between a pair of images $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ is computed as the average Euclidean distance between $\hat{\mathbf{p}}_{i}$ and $\mathbf{p}_{j}$. To analyze pose diversity under different confidence thresholds $\tau$, we compare the pose diversity scores of various methods across a range of $\tau$ values. As shown in the Fig. 8, our CoDi demonstrates a clear advantage over other SCG methods across all $\tau$ settings. We use $\tau=0.7$ in our experiments to balance keypoint reliability and coverage.
@@ Appendix C Limitations
Similar to prior subject-mask-based methods such as ConsiStory Tewel et al. (2024), our CoDi  framework relies on cross-attention scores to extract subject masks and estimate image token importance in the OT Plan. Occasionally, the pre-trained diffusion model assigns higher attention to background regions than to the subject, as shown in Fig. 9, which hinders the effective transport of identity features $\mathbf{X}_{\mathsf{id}}$ from the reference image $\mathbf{x}_{\mathsf{id}}$ to target image $\mathbf{x}_{n}$, resulting in subject inconsistency. However,such failures are rare in practice (under 5%) and can be solved by simply changing the seed.
Appendix D Inference Time and Memory Usage
Table 4: Inference Time and Memory Usage. We report the inference time (in seconds) and peak GPU memory usage on a single A6000 GPU for generating a set of five images from a prompt set with a resolution of $1024\times 1024$. StoryDiffusion Zhou et al. (2024) is excluded due to excessive GPU memory consumption beyond the A6000’s limit.
Method Inference time (s) GPU memory (GB)
Vanilla SDXL Podell et al. (2023) 77.67 35.58
1Prompt1Story Liu et al. (2025) 115.85 17.13
ConsiStory Tewel et al. (2024) 113.88 46.60
CoDi (ours) 154.89 45.20
We measure the inference time and memory usage of different SCG methods on a single A6000 GPU, as shown in Table 4. We report the wall-clock time for generating a set of five images from a prompt set (since the baseline method ConsiStory Tewel et al. (2024) performs cross-image attention across a batch of images) at a resolution of $1024\times 1024$. Based on Table 4, our method CoDi exhibits slightly higher inference time (154.89s) and comparable GPU memory usage (45.20GB) relative to ConsiStory. While 1Prompt1Story Liu et al. (2025) is the most memory-efficient, it compromises subject consistency. Note that we exclude StoryDiffusion Zhou et al. (2024) due to excessive GPU memory usage beyond the A6000’s limit at $1024\times 1024$ resolution (its original setting uses $768\times 768$).
Figure 10: CoDi shows superior results compared to SOTA training models.
Appendix E Comparison with SOTA Training Method
We compare CoDi with UNO Wu et al. (2025), as shown in Fig. 10. While UNO achieves subject consistency, its identical layout across outputs shows poor action prompt adherence and aesthetics. In contrast, CoDi demonstrates improved performance, highlighting the effectiveness of our training-free approach, which shows competitive results compared to SOTA company-level models.
Figure 11: CoDi (Cosine) fails to transport the corresponding features when there is a significant difference between the reference and target images. In contrast, CoDi (OT) achieves overall structural alignment, thereby preserving pose diversity.
Appendix F The Effectiveness of Optimal Transport
To assess the effectiveness of OT, we compare it with a simpler cosine similarity approach. CoDi (OT) achieves 0.0758, significantly outperforming CoDi (Cosine) at 0.0704 in pose diversity. As shown in Fig. 11, when there is a large feature difference between the reference and target images, especially across styles, CoDi (Cosine) tends to transport the mismatched features, failing to preserve the reference image’s pose. In contrast, CoDi (OT) effectively maintains the pose while enabling diverse variations, resulting in a failure to preserve the reference image’s pose. In contrast, CoDi (OT) effectively preserves the pose of the reference image while maintaining diverse pose variations.
Figure 12: CoDi can maintain both subject and clothing consistency while preserving pose diversity by adding the clothing description in the prompt and adjusting the subject mask threshold to include clothing in the foreground mask.
Appendix G Generalization to DiT-based Architectures
Figure 13: We have extended CoDi to the DiT-based architecture (Flux), with the generated results maintaining subject consistency while achieving diverse pose diversity.
We adapt CoDi to DiT-based models (Flux), with the generation results shown in Fig. 13, demonstrating richer details, subject consistency, and pose diversity.
Appendix H User Study Details
We conducted a user study comparing our method with state-of-the-art approaches, including ConsiStory, StoryDiffusion, and 1Prompt1Story. Thirty prompt sets, each containing four fixed-length prompts, were randomly sampled. Thirty-nine participants evaluated which method achieved the best overall image quality in terms of subject consistency, pose/layout diversity, and prompt fidelity.
Participants were instructed to select the set that best satisfied three evaluation criteria: subject consistency, pose/layout diversity, and prompt fidelity. Fig. 14 illustrates these criteria at the start of the questionnaire. To facilitate informed selections, a representative example was provided, accompanied by a performance comparison and rationale, highlighting the reasoning behind choosing the optimal set.
Figure 14: Questionnaire of user study. Evaluation Criteria outlines the standards for selecting image sets. Example illustrates a representative question for demonstration purposes. Performance Comparison and Selection Rationale demonstrates how the best choice is determined based on the example. These sections visually convey the evaluation criteria and guide participants’ selections.
Appendix I Additional Results
@@
Appendix J The Use of Large Language Models (LLMs)
In our work, large language models (LLMs) were employed primarily for general writing assistance. Specifically, we used LLMs to refine sentence expressions, check for grammatical errors, and convert tables into format.